• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fatiha

    Log entry time 00:08:19 on August 15, 2006

    Entry number 180014

    keyword=Shift Summary

     
    Shift crew: F. Benmokhtar, Y. Ilieva, X. Qian 
    

    -4:20pm P0 Left set to -5:25pm Beam went off, MCC is saying that there is a safety system trip

    that they need to investigate this. -6:10pm: Called MCC for an update, there is an RF problem in teh south Linac. This will take a while to repair. -6:50pm: Beam back -9:40pm: Inserted S0, downloaded the S0 trigger, increaded the cuurent to 10mu, set Ps4 =17, Ps3=20 and Ps6=5 -11:20pm: Switched to P0=0.4666 GeV/c

    Run Summary:

    Kin23b: P0 = 0.47615 GeV/c, current = 5 uA 2716 LH2 1.9 M .... field was a bit lower than what it should be. 2717 LH2 1.9 M good run 2718 LH2 3.6 K good run, beam went off 2719 LH2 1.9 M 2720 LH2 1.9 M 2721 LH2 1.9 M 2722 LH2 1.9 M 2723 LH2 1.9 M 2724 LH2 1.9 M 2725 LH2 1.9 M 2726 LH2 1.9 M 2727 LH2 1.9 M 2728 LH2 1.9 M 2729 LH2 1.9 M

    Kin23d: Elastic Cross section measuremnts; 2732 JUNK test 2733 JUNK test 2734 JUNK test 2735 LH2 1.9 M current = 10 uA, Deadtime = 7% 2736 LH2 1 M current = 5 uA, Deadtime = 3%

    Kin24: P0=0.4666 GeV/c 2738 LH2 1.9 M 2739 LH2 1.9 M 2740 LH2 goes to the next shift.